∫ sin³ x cos³ x dx


1
Expert's answer
2021-04-15T07:19:46-0400

"\int \sin^3 x \cos ^3 x dx"

Let "t =\cos x" , "dt = - \sin x dx" . Then "\int \sin^3 x \cos ^3 x dx = \int -\sin^2x \, t^3 dt"

Also, "\sin^2x = 1 -\cos^2x = 1-t^2" . So integral can be re-written as

"\displaystyle \int -(1-t^2) t^3 dt = - \int(t^3-t^5)dt = \frac{-t^4}{4} + \frac{t^6}{6} + C = \frac{\cos^6x}{6} - \frac{\cos^4x}{4}+C"


Answer: "\displaystyle \frac{\cos^6x}{6} - \frac{\cos^4x}{4}+C"